OpenStudy (anonymous):

A circle is intercepted by five central angles each measures m, n, (p + 45), 75, and 111 degrees. Which is m + n + p?

OpenStudy (anonymous):

m +n + p + 45 = 360 -75 -111 (total at centre = 360) m + n +p = 360 -75 -111 -45 =129 degrees

OpenStudy (anonymous):

Thankyou so much!!!!!!![:

OpenStudy (anonymous):

u r welcome